Hockey teams receive 2 points when they win and 1 point when they tie. One season, a team won a champuionship with 60 points. They won 12 more games than they tied. How many wins and how many teams did they have?
Answer by htmentor(1343)

Let x = the number of ties
Then x + 12 = the number of wins
Then the number of points is
2(x + 12) + x = 60
3x + 24 = 60
3x = 36
x = 12
Therefore, the number of wins was 24, the number of ties was 12
Check:
24*2 + 12*1 = 48 + 12 = 60